OpenStudy (i_love_my_nieces):

a $4,000 principal earns 5% interest, compounded annually. After 4 years, what is the balance in the account? A $3,500 principal earns 3% interest, compound semiannually. After 20 years, what is the balance in the account? A tractor costs $15,450 and depreciates in value by 14% per year. How much will the tractor be worth after 3 years?

OpenStudy (i_love_my_nieces):

@mathstudent55

OpenStudy (i_love_my_nieces):

@Directrix

OpenStudy (i_love_my_nieces):

@dan815

OpenStudy (i_love_my_nieces):

@skullpatrol

OpenStudy (i_love_my_nieces):

@Teddyiswatshecallsme

OpenStudy (i_love_my_nieces):

@whpalmer4

Directrix (directrix):

I used an online calculator. See what you think about it. http://www.thecalculatorsite.com/finance/calculators/compoundinterestcalculator.php

Directrix (directrix):

Did you get this?

OpenStudy (i_love_my_nieces):

Yeah. What about the 2nd and 3rd?

Directrix (directrix):

If you use the online calculator, you can knock these out quickly. I'll check answers with you. http://www.thecalculatorsite.com/finance/calculators/compoundinterestcalculator.php

OpenStudy (i_love_my_nieces):

Can you do them quickly please this has to be turned in at 11

Directrix (directrix):

Go to the calculator site and enter the data as shown on the attachment. Punch calculate. You'll be finished well before 11 if you do that.

OpenStudy (i_love_my_nieces):

I did and it is not working my computer is a P.O.C

Directrix (directrix):

Then, you can try using the formula with your calculator.

OpenStudy (i_love_my_nieces):

1. 4,862.03 2. 6,349.06 3. 22,889.85 Are these correct?

OpenStudy (i_love_my_nieces):

Are they????

Directrix (directrix):

>>3. 22,889.85 cannot be correct because the tractor would be worth more after 3 years of depreciation than it was worth at the start.

OpenStudy (i_love_my_nieces):

9,827.07?

Directrix (directrix):

15450*(.86)^3= ?

OpenStudy (i_love_my_nieces):

9,827.07

Directrix (directrix):

>9,827.07? That is what I got.

Directrix (directrix):

1. 4,862.03 Yes

OpenStudy (i_love_my_nieces):

One more and I would like work shown one how to do it, please

Directrix (directrix):

Use the compound interest formula. P = 4000 * (1 + .04/1)^3 P = What is this equal to: 4000 * (1 + .04/1)^3

Directrix (directrix):

Give answer to the nearest dollar.

OpenStudy (i_love_my_nieces):

4499.456

OpenStudy (i_love_my_nieces):

4499.46?

Directrix (directrix):

Yes $4,499.46 to the nearest dollar is 4,499.00

OpenStudy (i_love_my_nieces):

So can you put all of the step into one post to I can see it all please thank you

Directrix (directrix):

P = C* (1 + r/n)^(n*t) P = 4000 * (1 + .04/1)^3 P =